OpenStudy (anonymous):

as the shuttle bus comes to a sudden stop to avoid hitting a dog, it accelerates uniformly at -4.1 m/s2 as it slows from 9.0 m/s to 0 m/s. find the time interval of aceleration for the bus.

OpenStudy (amistre64):

\[a=\frac{2d}{t^2}\]

OpenStudy (anonymous):

nt exactly a question for math forum use speed time equations v=u+at s=ut+0.5at^2 or v^2=u^2-2as

OpenStudy (amistre64):

\[v^2=(v_o)^2+2ad\] still not quite right lol \[v=v_o+at\]

OpenStudy (amistre64):

therefore:\[t=\frac{v_f-v_i}{a}\]

OpenStudy (anonymous):

Im still confused..